OpenStudy (jakyfraze):

Solve the following inequality: 38 < 4x + 3 + 7 – 3x. A. x < 28 B. x > 28 C. x < 4 D. x > 4

OpenStudy (jakyfraze):

@breanna1999 does this follow the same rule like the other one?

OpenStudy (breanna1999):

This one does but has a few added steps

OpenStudy (jakyfraze):

@breanna1999 like what?

OpenStudy (breanna1999):

You must first change the < sign to an equal sign so you can solve the problem

OpenStudy (jakyfraze):

@breanna1999 how? and what do you mean

OpenStudy (breanna1999):

The right side of the problem has your terms that you must combine and the left side a single term all by itself. This problem is asking you to solve it so you will have 38<4x+3+7-3x becomes 38=4x+3+7-3x

OpenStudy (jakyfraze):

@breanna1999 ohh okay, so after that whats next?

OpenStudy (breanna1999):

Combine your like terms on the right side

OpenStudy (jakyfraze):

@breanna1999 starting with 4x + -3x ?

OpenStudy (breanna1999):

Yes and 4x+(-3x)=what's

OpenStudy (jakyfraze):

1x ?

OpenStudy (breanna1999):

Yes or you could just say x. Now to the next one

OpenStudy (jakyfraze):

why just say x when it's x1 ?

OpenStudy (breanna1999):

1x is equal to x because there is an imaginary 1 in front of the x

OpenStudy (jakyfraze):

how?

OpenStudy (jakyfraze):

if it was 2x 3x 4x 5x 6x it will still equal x?

OpenStudy (breanna1999):

You can just write it as 1x If it helps you

OpenStudy (jakyfraze):

I'm just wondering

OpenStudy (breanna1999):

My teacher always taught me that if you have 1x you can just leave it as x because of the understood 1 in front of it

OpenStudy (jakyfraze):

so even if it was 2x and up ? it will still be as x. or is it just the number 1x = x

OpenStudy (breanna1999):

No it's only if you see 1x

OpenStudy (jakyfraze):

so your saying if it was 2x it will equal 2x, right?

OpenStudy (jakyfraze):

sorry if I'm confusing you, I'm just confuse my self :(

OpenStudy (breanna1999):

Yes 2x=2x and 3x=3x and so on

OpenStudy (jakyfraze):

ohh okay thank i was confuse for a minute. so whats the difference between the first question i ask and this one?

OpenStudy (jakyfraze):

Simplify this expression: 4p + 9 + (–7p) + 2 = ? A. 3p + 7 B. –3p + 11 C. 3p + 11 D. 11p + 11

OpenStudy (breanna1999):

This one involes you needing to solve it

OpenStudy (jakyfraze):

Solve the following inequality: 38 < 4x + 3 + 7 – 3x. A. x < 28 B. x > 28 C. x < 4 D. x > 4

OpenStudy (breanna1999):

So go ahead and combine your like terms on this equation

OpenStudy (jakyfraze):

ohh so the first question i did not solve it?

OpenStudy (jakyfraze):

but this one i have to solve it?

OpenStudy (breanna1999):

You solved by combining like terms for the first one for the second one you have a two step equation

OpenStudy (breanna1999):

So far we have 38=1x+3+7

OpenStudy (jakyfraze):

38=x 10?

OpenStudy (breanna1999):

38=x+10 is correct now we have to solve the equation. Do you know how to do this?

OpenStudy (jakyfraze):

im not really sure, can you show me please?

OpenStudy (breanna1999):

Yes I can 38=x+10 38-10=x+10-10 28=x 28<x

OpenStudy (jakyfraze):

i dont understand the second step. how did you do that or get that? I'm sorry :(

OpenStudy (breanna1999):

This question is askin you to solve to find the inequality. You must have x on one side by itself. In order to do this you must subtract 10 from both sides

OpenStudy (jakyfraze):

so subtract 10 from 38 and 10?

OpenStudy (breanna1999):

Yes 10-10 equals 0 and 38-10 equals 28

OpenStudy (jakyfraze):

ohh, wow. were did you learn this? i want to learn more of this

OpenStudy (breanna1999):

I learned it in algebra 1. There's a website called Khan Academy that can go into further detail

OpenStudy (jakyfraze):

@breanna1999 is it free?

OpenStudy (breanna1999):

It is

OpenStudy (jakyfraze):

okay thank you

OpenStudy (jakyfraze):

so the answer is B. right?

OpenStudy (breanna1999):

Yes

OpenStudy (jakyfraze):

yessss! thank you